Find the integers. Write them in numerical order from least to greatest, separated with commas and NO spaces.
The sum of four consecutive integers is -22.

Answer by stanbon(75887)

Find the integers. Write them in numerical order from least to greatest, separated with commas and NO spaces.
The sum of four consecutive integers is -22.
Let the integers be:
x-2,x-1,x,x+1
----
Equation:
x-2 + x-1 + x + x+1 = -22
4x-2 = -22
4x = -20
x = -5
----
x-2 = -7
x-1 = -6
x = -5
x+1 = -4
-----
Ans: -7,-6,-5,-4
--------------
Cheers,
Stan H.
==============